Emergency relief supplies was dropped from a helicopter flying at a height of 7.5 m while traveling at 15 m/s. Calculate the horizontal distance traveled by the package before it hits the ground.
Answer by Theo(13342)

the height that the package was dropped at is 7.5 meters.

the helicopter is traveling at a horizontal speed of 15 meters per second.

the distance a package dropped in free fall is modeled by the formula:

d = 1/2 * 9.80665 * s^2.

d is the distance
s is the number of seconds.
9.80665 * s^2 is the free fall acceleration due to gravity.

when d = 7.5 meters, the formula becomes:

7.5 = 1/2 * 9.80665 * s^2

multiply both sides of the equation by 2 and divide both sides of the equation by 9.80665 to get:

7.5 * 2 / 9.80665 = s^2

solve for s^2 to get:

s^2 = 7.5 * 2 / 9.80665 = 1.529574319.

take the square root of both sides of the equation to get:

s = sqrt(1.529574319) = 1.236759605.

that's how long it takes the package to hit the ground from a height of 7.5 meters in free fall.

since the helicopter is traveling a horizontal distance at the constant rate of 15 meters per second, then the package will have traveled a horizontal distance of 15 meters per second * 1.236759605 seconds = 18.55139407 meters when it hits the grouond.

if i did this correctly, 18.55139487 meters should be your answer.